# LEMON-Mapping: Loop-Enhanced Large-Scale Multi-Session Point Cloud Merging and Optimization for Globally Consistent Mapping

### v1

@@ IV-C Multi-Robot Study
Figure 8: The map merging result of the S3E library. The local maps generated by three methods are selected for comparison. Our method almost eliminates the divergence of local multi-session point cloud maps. While LAMM [29] and DCL-SLAM [27] suffer from serious inconsistencies.
Figure 9: The optimized trajectory, position, and position error for proposed method in Island ((a)-(c)) and Town ((d)-(f)).
@@
We select a scene where all three methods successfully merge the multi-robot point cloud for display. Figure 8 visualizes the results from the S3E Library sequence. The left image shows our globally consistent and accurate merged map, while the right part shows the local map fusion of the three methods and the comparison of their respective trajectories. Our method yields tightly aligned local maps, whereas LAMM and DCL-SLAM exhibit severe local divergences. The comparison of trajectory and ground truth further confirms the superior accuracy of our method. Fig. 10 and Fig. 11 present multi-robot trajectories comparisons between our method, LAMM, and DCL-SLAM. The trajectories estimated by our method closely follow the ground truth for each robot, exhibiting low and stable errors. In contrast, the benchmark methods show significant deviations from the ground truth in some regions and minor local inconsistencies in others. The large deviations are likely caused by incorrect loop closures, while the local inconsistencies may result from the lack of bundle adjustment, which fails to eliminate the divergence of submaps.
The performance disparity can be attributed to LAMM and DCL-SLAM relying solely on PGO with loop closure constraints. Although this achieves general global consistency, it often neglects the refinement of overlapping local regions. Lack of attention to the geometric structure of point cloud maps leads to serious local multi-robot map divergence. Our framework explicitly re-examines the role of loop and enhances its utilization by performing spatial bundle adjustment in local regions, while propagating the refined results globally via the last pose graph optimization. This process effectively reduces divergence and ensures consistent multi-session map fusion. Figure 9 shows optimized trajectories and error plots for different scenes, the multi-robot trajectories generated by our framework are basically consistent with the ground truth, and the error is stable and small, further demonstrating the superior performance and stability of our method.
TABLE V: LOOP CLOSURE RECALL DATA
Scene Initial After Rejection After Recalling RMSE $w/o$ LR RMSE $w/$ LR
@@
IV-D Ablation Study
TABLE VI: ABLATION STUDY RESULTS
         Scene FPGO FPGO + BA Lemon Full
@@ IV-D Ablation Study
Figure 10: Comparison of multi-robot trajectories and error plots between our method, LAMM, and DCL-SLAM on the S3E Library dataset. The ground truth is shown as a dashed line in the figure.
Figure 11: Comparison of multi-robot trajectories and error plots between our method, LAMM, and DCL-SLAM on the S3E Tunnel dataset. The ground truth is shown as a dashed line in the figure.
Figure 12: Result of ablation study. (a) and (b) show the merging map of FPGO and FPGO with BA. (c) and (d) show the merging map of FPGO with BA and Lemon mapping full model. The comparison regions are framed in red.
Figure 13: The map merging result of five sessions in the R3LIVE HKU Park dataset. The local maps of (a)-(f) are enlarged to show the details.
To understand the contribution of each component in our framework, we perform ablation experiments on the S3E, GEODE, and MARS-LVIG datasets. Our ablation experiment consists of two parts.
First, we conduct an ablation experiment on the loop recall module in III-B. We compare the performance of the first pose graph optimization with and without loop recall (LR) using RMSE of the ATE (m), reported in the table as “RMSE $w/$ LR” and “RMSE $w/o$ LR,” respectively. Because the S3E dataset contains a large number of complex scenes that combine unstructured and structured data, as well as many scenes with high similarity, it is easy to have a large number of false loops, so we use it as the ablation dataset for loop recall.
Table V shows the number of loops in different loop processing stages, as well as the results of the first pose graph optimization with and without loop recall. The experimental results show that the first PGO achieves smaller errors and better performances when the mistakenly rejected loops are recalled. Beyond its impact on PGO, loop recall also introduces additional constraints that benefit subsequent bundle adjustment.
@@
Table VI summarizes the results, with the best values highlighted in bold. The complete framework consistently achieves the lowest RMSE, validating the effectiveness of combining local BA and two PGO steps. For the variant of FPGO + BA, although local BA improves relative accuracy, it may disrupt the continuity of odometry since it only refines loop regions. Consequently, its standalone use can increase RMSE compared to FPGO alone. However, the final PGO integrates both local BA constraints and odometry continuity, propagating local accuracy to the global map and significantly reducing RMSE.
Fig. 12 shows that merging maps of three variants above. (a) and (b) demonstrate the function of spatial bundle adjustment to improve local consistency and accuracy compared to PGO only. (c) and (d) in Fig. 12 further illustrates that maps refined with both BA and last pose graph optimization have better global consistency than using BA alone, confirming the role of last PGO in maintaining global map structure.
IV-E Scalability Study
@@
       Dataset Robot Num Merged Num Success Rate
@@
      HKU Mainbuilding 20 18 90%
Most existing map fusion frameworks are limited to scenarios involving only a few robots (less than three), and their scalability to large-scale deployments remains unproven. To assess the scalability of our proposed framework, we conduct experiments on the R3LIVE dataset, which we subdivide into groups of 5, 10, and 20 sessions. A successful fusion is defined as the correct alignment of each session with all its adjacent sessions that shares sufficient map overlap.
Table VII summarizes the corresponding results of three cases. Our framework achieves a 100% success rate in the 5-session and 10-session scenarios. In the more challenging 20-session case, a single failure occurred since its associated loop is not detected, resulting in a success rate of 90%. These results demonstrate the scalability and robustness of the Lemon-Mapping system when dealing with large numbers of multi-robot map merging.
Fig. 13 visualizes the fused map for the 5-session case of HKU Park in R3LIVE dataset, including a global point cloud map from a bird’s eye view (BEV) and 6 enlarged map details. The five-session map shows good fusion effect and local accuracy in both structured and unstructured environments.
V Conclusion
This paper presents LEMON-Mapping, a Loop-Enhanced Large-Scale Multi-Session Point Cloud Merging and OptimizatioN framework for Globally Consistent Mapping. Lemon-Mapping is a scalable framework for large number of robots. Unlike existing methods that rely solely on pose sequences, it enhances the role of loop closures by directly utilizing geometric constraints from overlapping point clouds and reasonably transferring the local accuracy to the global map. The framework includes a loop processing module for reliable loop selection and innovate loop recall, and a map merging module that combines two-step pose graph optimization with a spatial bundle adjustment tailored for multi-robot systems. The spatial bundle adjustment operates on loop-based spatial windows, allowing geometry-aware pose optimization across unordered and cross-robot trajectories. To ensure global consistency, we sparsify the BA constraints and propagate local accuracy across the entire map using a refined pose graph optimization strategy. Experimental results on S3E, MARS-LVIG, GEODE, R3LIVE and our self-collected dataset demonstrate that our approach maintains high success rates and robustness as the number of sessions increases, while outperforming traditional methods in geometric accuracy and scalability.
@@ V Conclusion
A standard result on the continuous dependence of polynomial roots on coefficients now implies that each ordered root $\lambda_{i}(x)$ varies continuously with the coefficients $\{c_{k}(x)\}$, and hence with $x$ itself. Therefore every eigenvalue function $x\mapsto\lambda_{i}(x)$ is continuous.

### v2

@@ VII-C Multi-Robot Study
We select a scene where all three methods successfully merge the multi-robot point cloud for display. Fig. 8 visualizes the results from the S3E Library sequence. The left image shows our globally consistent and accurate merged map, while the right part shows the local map fusion of the three methods and the comparison of their respective trajectories. Our method yields tightly aligned local maps, whereas LAMM and DCL-SLAM exhibit severe local divergences.
The comparison of trajectory and ground truth further confirms the superior accuracy of our method. Fig. 9 shows optimized trajectories and error plots for different scenes, the multi-robot trajectories generated by our framework are basically consistent with the ground truth, and the error is stable and small, further demonstrating the superior performance and stability of our method. Fig. 10 and Fig. 11 present multi-robot trajectories comparisons between our method, LAMM, and DCL-SLAM. The trajectories estimated by our method closely follow the ground truth for each robot, exhibiting low and stable errors. In contrast, the benchmark methods show significant deviations from the ground truth in some regions and minor local inconsistencies in others. The large deviations are likely caused by incorrect loop closures, while the local inconsistencies may result from the lack of bundle adjustment, which fails to eliminate the divergence of submaps.
The performance disparity can be attributed to LAMM and DCL-SLAM relying solely on PGO with loop closure constraints. Although this achieves general global consistency, it often neglects the refinement of overlapping local regions. Lack of attention to the geometric structure of point cloud maps leads to serious local multi-robot map divergence. Our framework explicitly re-examines the role of loop and enhances its utilization by performing spatial BA in local regions, while propagating the refined results globally via the last pose graph optimization. This process effectively reduces divergence and ensures consistent multi-session map fusion.
VII-D Ablation Study
TABLE V: LOOP CLOSURE RECALL COMPARISON
Sequence Initial After Rejection After Recalling RMSE $w/o$ LR RMSE $w/$ LR
@@ VII-D Ablation Study
TABLE VI: RMSE OF THE ATE(m) OF ABLATION STUDY
         Sequence FPGO FPGO + BA LEMON Full
@@
Figure 12: (a) shows the global map in Laboratory with loop recall. (b) and (c) present enlarged views of the same region without and with loop closure, respectively. In (c), the recalled loop information contributes to improved alignment and consistency across multi-robot maps.
Figure 13: Result of ablation study in our Flying Arena dataset. (a) and (b) show the merged map of FPGO and FPGO + BA. (c) and (d) show the merged map of FPGO + BA and LEMON-mapping full model. The comparison regions are framed in red.
Figure 14: The map merging result of the five-session dataset in R3LIVE HKU Park. The local maps of (a)-(f) are enlarged to show the details.
To understand the contribution of each component in our framework, we perform ablation experiments on the S3E, GEODE, MARS-LVIG datasets, as well as our self-collected dataset. All sequences of our self-collected dataset are divided into two sessions. Our ablation experiment consists of the following two parts.
First, we conduct an ablation experiment on the loop recall in Section IV. We compare the performance of the first pose graph optimization with and without loop recall (LR) using RMSE of the ATE (m), reported in the table as “RMSE $w/$ LR” and “RMSE $w/o$ LR,” respectively. Because the S3E dataset contains a large number of complex scenes that combine unstructured and structured data, as well as many scenes with high similarity, it is easy to have a large number of false loops, so we use it as the ablation dataset for loop recall.
Table V shows the number of loops in different loop processing stages, as well as the results of the first pose graph optimization with and without loop recall. The experimental results show that the first PGO achieves smaller errors and better performances when mistakenly rejected loops are recalled. Beyond its impact on PGO, loop recall also introduces additional constraints that benefit subsequent bundle adjustment.
Fig. 12 illustrates the results of the first ablation experiment in our laboratory scene. (a) presents the global map generated with the loop recall step. (b) and (c) compare the local mapping results without and with loop recall, respectively. It is evident that the loop recall introduces additional constraints that significantly improve point cloud alignment within the selected region.
@@
Table VI summarizes the results, with the best values highlighted in bold. The complete framework consistently achieves the lowest RMSE, validating the effectiveness of combining spatial BA and two PGO steps. For the variant of FPGO + BA, although local BA improves relative accuracy, it may disrupt the continuity of odometry since it only refines loop regions. Consequently, its standalone use can increase RMSE compared to FPGO alone. However, the last PGO integrates both local BA constraints and odometry continuity, propagating local accuracy to the global map and significantly reducing RMSE.
Fig. 13 shows the map merging results of three variants above in our Flying Arena scene. (a) and (b) demonstrate the function of spatial BA to improve local consistency and accuracy compared to PGO only. (c) and (d) in Fig. 13 further illustrates that maps refined with both BA and last pose graph optimization (LPGO) have better global consistency than using BA alone, confirming the role of LPGO in maintaining global map structure.
VII-E Scalability Study
@@
       Sequence Robot Number Merged Number Success Rate
@@
      HKUST Campus 20 20 100%
Most existing map fusion frameworks are limited to scenarios that involve only a few robots (less than five [29, 38, 27, 28]), and their scalability to large-scale deployments is not yet proven. To assess the scalability of our proposed framework, we perform experiments on the R3LIVE dataset, which we subdivide into groups of 5, 10, and 20 sessions. A successful fusion is defined as the correct alignment of each session with all its adjacent sessions that shares sufficient map overlap.
Table VII summarizes the corresponding results of three cases. Our framework achieves a success rate 100% in all experiments, including the five-session, ten-session, and twenty-session scenarios. These results demonstrate the scalability and robustness of the LEMON-Mapping system when dealing with large numbers of multi-robot map merging.
Fig. 14 visualizes the fused map for the five-session case of HKU Park in R3LIVE dataset, including a global point cloud map from a bird’s eye view (BEV) and 6 enlarged map details. The five-session map shows good fusion effect and local accuracy in both structured and unstructured environments.
VIII Conclusion
This paper presents LEMON-Mapping, a Loop-Enhanced Large-Scale Multi-Session Point Cloud Map Merging and OptimizatioN framework for Globally Consistent Mapping. LEMON-Mapping is a scalable framework for a large number of robots. Unlike existing methods that rely solely on pose sequences, it enhances the role of loop closures by directly utilizing geometric constraints from overlapping point clouds and reasonably transferring the local accuracy to the global map. The framework includes a loop processing module for reliable loop selection and innovate loop recall, and a map merging module that combines two-step pose graph optimization with a spatial bundle adjustment tailored for multi-robot systems. The spatial bundle adjustment operates on loop-based spatial windows, allowing geometry-aware pose optimization across unordered and cross-robot trajectories. To ensure global consistency, we sparsify the BA constraints and propagate local accuracy across the entire map using a refined pose graph optimization strategy. Experimental results in S3E, MARS-LVIG, GEODE, R3LIVE and our self-collected dataset demonstrate that our approach maintains high map merging success rates and robustness as the number of sessions increases, while outperforming traditional methods in geometric accuracy and scalability.
@@ VIII Conclusion
A standard result on the continuous dependence of polynomial roots on coefficients now implies that each ordered root $\lambda_{i}(x)$ varies continuously with the coefficients $\{c_{k}(x)\}$, and hence with $x$ itself. Therefore, every eigenvalue function $x\mapsto\lambda_{i}(x)$ is continuous.